OpenStudy (anonymous):

What is the equation of the parabola, in vertex form, with vertex at (2,-4) and directrix y = -6? a. (y+4)^2=8(x-2) b. (y+6)^2=-8(x+2) c. (x+2)^2=8(y+4) d. (x-2)^2=8(y+4)

OpenStudy (experimentx):

must be d

OpenStudy (campbell_st):

since the vertex is y = -6 the parabola is in the form (y -k)^2 = 4a(x -h) the distance from the directrix to the vertex is 2 so a = 2... the parabola is concave right... (y + 4)^2 = 8(x - 2)

OpenStudy (anonymous):

is it A or D ?

OpenStudy (campbell_st):

its A

OpenStudy (anonymous):

its D :/

OpenStudy (experimentx):